OpenStudy (anonymous):

Find the circumference of point A. angle is 50 degrees and arc is 7.2. a) 103.68 b) 25.92 c) 48.9 d) 50.4 e) 51.84

OpenStudy (anonymous):

E

OpenStudy (divanshu):

e

OpenStudy (anonymous):

The EQUATION for calculating the ARC ANGLE RELATIONSHIP is theta/360° = s/C = s/2πr. So you have 50/360=7.2/C, C=51.84. Your answer is E.